OpenStudy (anonymous):

solve by the elimination method. 8x-9y=8.5, 7y-2x=-4.5

OpenStudy (mathstudent55):

Solve opne equation for one variable. Then substitute in the other equation. I suggest you start by solving the second equation for x.

OpenStudy (anonymous):

8x-9y=8.5 solve for x 8x = 8.5+9y divide both sides by 8 x = (8.5)/(8) + (9y)/8 substitute this value for x into the next equation to solve for y. you will be able to get a number answer for y. then plug the y number into either equation to get the number value for x.

OpenStudy (anonymous):

would it be 17.5/8?

OpenStudy (anonymous):